# Montréal, QC Canada

### v1

@@ 3 Motion Detection with DMD
Having now reviewed the necessary DMD-related methods, we present our method of motion detection. This application of DMD is the main contribution of this paper and validation of the method is left to the next section. Recall that a visual overview of the method was presented in Figure 2 previously.
@@
Input: $W\in\mathbb{R}^{M\times N+1}$, $r$, $p$, $T$
@@
The method is initiated by applying cDMD to windowed subsets of the streaming video data, which has been transformed into greyscale and vectorized into a matrix $W\in\mathbb{R}^{M\times N+1}$. The compression matrix $C$ and target rank $r$ are held constant over all windows, and each window is comprised of $T>0$ consecutive frames of video. Algorithm 1 provides our windowed cDMD method where we see that the output is the modulus of the continuous-time eigenvalues (5) within each window. The scope of this work is primarily limited to video data where the camera position remains fixed, meaning the background is assumed to be relatively unchanged over the frames. This means that if there is little or no motion in the video, the eigenvalues within the window will all correspond approximately to the background, i.e., $\omega_{i}\approx 0$. However, if motion is present in a window one expects to see this represented as at least one of the eigenvalues deviating from the background cluster. As one sweeps across windows, motion entering the frames is expected to be represented by a large spike in the DMD matrix spectrum, as shown in Figure 6 below.
Motion detection from the DMD spectrum is automated by considering the sets of eigenvalues $\{\omega_{i}^{(k)}\}_{i=1}^{r}$ and $\{\omega_{i}^{(k+1)}\}_{i=1}^{r}$ from the consecutive $k$th and $(k+1)$st windows. We compute the averages
@@
Then, defining a threshold $\Delta^{*}>0$, motion is flagged in the $(k+1)$st window if
@@
By tracking the relative change in the mean of the eigenvalues over windows we allow for the method to exclude background motion that persists over frames, such as leaves being rustled by the wind. Note that we need only save information from the $k$th and $(k+1)$st consecutive windows at a time, while if no motion is detected we discard information from the $k$th window and compare the spectrum of the $(k+1)$st and $(k+2)$nd windows. Combined with the data reductions from cDMD, this mean that only two $r\times r$ matrices are stored in memory to perform the method.
In practice, it was observed that foreground motion in video was sometimes accompanied by multiple consecutive activations; that is, the moduli of the eigenvalues would continue to increase in a few of the windows immediately following that of the initial detection. One way this can be resolved is by storing the (binary) results of the motion detection test for the previous few windows, and additionally checking that no motion was already flagged in these previous windows.
Finally, if motion is detected in a window, we apply the method of background separation from § 2.4. This allows for the isolation of the detected motion, again using only information from the stored $r\times r$ DMD matrix corresponding to the window where motion is first detected. This is summarized in the flowchart in Figure 4.
@@
This method involves initializing a number of parameters that may affect its performance. Precisely, the target rank $r$ dictates the number of eigenvalues in each window, while the target dimension $p$ from cDMD, the window length $T$, and most importantly, the detection threshold $\Delta^{*}$ all have bearing on the performance. The choice of parameters that leads to optimal performance of the method is problem-dependent and should be optimized with simple tests. This might include designing motion to be captured by the method, such as having someone walk into frame, while also decreasing the sensitivity to exclude frivolous motion that might be unique to the application. In the following section we demonstrate the method’s performance on real video data, while also evaluating performance with receiver operating characteristic curves and tuning the threshold detection parameter with $k$-fold cross-validation.
@@ 4 Experimental Validation
In this section we demonstrate the performance of our method on real video data. A full code repository to reproduce the results of this section is available. This includes all files used for our tests, as well as full MATLAB scripts to perform the method on video data and all of the coming validations: https://github.com/marco-mig/dmd-motion-detection.
In what follows, we fix the parameters $T$, $p$, and $r$, while focusing on the importance of the threshold parameter $\Delta^{*}$. Precisely, we take $p=20$ and $r=5$, meaning each DMD matrix is only $5\times 5$ dimensional. Thus, eigenvalues within each window are can be computed quickly without presenting a bottleneck to the real-time nature of the algorithm. Window sizes are fixed to be $T=80$, corresponding to 8/3 seconds. The random matrix $C$ is constructed only once and applied to all videos for consistency.
@@ 4.1 Dataset
Initial tests of the method featured free videos from the Background Models Challenge database BMC , including the one featured in Figure 1. However, to ensure a diversity of situations across the test set, we have opted to film our own videos using the built-in camera on the iPhone 13. Our videos were filmed to simulate security footage in that the camera is held still and various individuals enter and exit the frame to be detected. These videos include numerous variations such as the speed of entry and/or exit of the individual, the distance of the motion from the camera, the time the individual spends in the frame, and the actions performed by the individuals in the frame. Our test set includes a total of 20 videos, of which 15 were filmed in daylight and 5 with indoor lighting. Videos are of size 426 by 240 pixels and range in length from 27 seconds (810 frames) to 109 seconds (3270 frames). Each video begins and ends with a few seconds of a still background shot to provide a baseline for the method to gauge performance. Frames from videos in two lighting conditions are provided in Figure 5.
@@ 4.2 Performance Demonstration
Our experiments demonstrate that sudden motion from an individual entering or exiting the video frame is accompanied by a spike in the average eigenvalues, as measured by (16). A characteristic demonstration is provided in Figure 6 where one can clearly see the motion detection via large deviations in the DMD spectrum. We have observed that whenever an individual enters the frame, it is reflected in the eigenvalues at the window number corresponding closely to the frame number minus the window length, as this is when the frame at which the individual enters first appears in the window as the rightmost column of the data matrix. However, when individuals exit the frame, the window number where the change in eigenvalues is observed corresponds to the frame number itself. This is because it is the last window in which information about the movement is present in the data matrix, and so the change occurs after this information (in the leftmost column of the data matrix) vanishes in the subsequent window. To summarize, there is a temporal delay of length $T$, the window size, in detecting when an object has left the frame. Thus, a distinction is made between objects entering and exiting the frame, and the frame numbers which are supplied to the algorithm which calculates the error score in the following subsection have been adjusted accordingly.
@@ 4.3 Performance Evaluation
To display the relationship between the true positive rate (TPR) and false positive rate (FPR) we will use a receiver operating characteristic (ROC) curve (see, e.g., stat ). The TPR can be thought of the rate at which the algorithm successfully classifies positive events as positive, while the FPR is the rate at which the algorithm classifies negative events as positive. A perfect classifier necessarily has a TPR of 1 and an FPR of 0, and the ROC curve for such a classifier is the graph of the indicator function of the interval $(0,1]$. Our ROC curve is traced out by recording the TPR and FPR of the algorithm over threshold values $\Delta^{*}$ ranging from $10^{-10}$ to $10^{10}$. In this section, we do not implement the fix for consecutive activations as described in § 3 as this interferes with the expected behaviour of the algorithm at threshold values close to zero. That is, for these values, activations at nearly every (if not every) window containing motion is expected. Finally, a half-second tolerance (forward and backward) in detecting each event is imposed.
@@ 4.4 Optimizing the Threshold
While our method contains a number of parameters to be set by the user, none is as important as the threshold $\Delta^{*}$. In this subsection we present a method of determining an optimal threshold to get the best performance from the method using a database of test videos that should be curated by the user before applying the method for video surveillance. We begin by defining an appropriate error measure which can then be minimized by the optimal threshold.
@@ 4.4.1 Measuring Error
Recall from (§ 4.3) that there are a total of four possible outcomes of the algorithm: either it classifies correctly (encompassing TPs and TNs) or incorrectly (FPs and FNs). In many applications, these two latter outcomes are not equally as costly. This is because in the context of video surveillance, a FN, such as not detecting an intruder, can be much more costly than a FP, such as falsely indicating the presence of an intruder. Thus, our goal here is to present a possible error metric that takes this into consideration.
@@ 4.4.2 A Variant of k-Fold Cross-Validation
We will use a variant of $k$-fold cross-validation (see, e.g., stat ) to determine an optimal threshold value using our error metric (17) and database of videos. In what follows we fix $d^{*}=30$, which corresponds to a tolerance of detection within 1 second of an event taking place. The penalty ratio in the error metric is taken to be $c=100$.
@@
Error $E_{i}$ 67.8 72.2 49 170.8
We implement the above with $k=4$ since our dataset is relatively small, and we initialize $P$ with 1000 evenly-spaced values between 0 and 1. Table 1 presents our results for one application of the modified $k$-fold cross-validation procedure. It can be seen that the optimal threshold corresponds to an error score of 49, where we recall that FP and FN have a score of 1 and 100, respectively. However, due to the small size of the video databases, there are large discrepancies between the error scores corresponding to the optimal thresholds. Moreover, each possible arrangement of the database into the training and validation sets results in different values. We expect that this cross-validation method will produce a more representative set of optimal thresholds and corresponding errors for a larger database.
@@
From these results, two observations can be made. First, while the eigenvalue detection system can be effective when optimized for a single video, as shown in Figure 6, when optimizing over a set of videos, the performance tends to decrease. In particular, more FPs are made in order to minimize the occurrence of FNs over all the videos in the training set. This is evidently a consequence of the weight given to FN errors, and is thus dependent on the application. The second observation is that the optimal thresholds and their corresponding errors are always a trade-off between FPs and FNs, since they vary heavily depending on how each type of error is weighed according to the value of $c$. As demonstrated in Figure 8, plotting the average error over each of the training sets on which the cross-validation procedure is applied results in a curve with a distinct global minimum. Large error is observed for very small and very large threshold values due to the abundance of FP and FN errors, respectively.
@@ 5 Conclusion
The method relies on spikes in the DMD spectrum coming from sudden motion in the frame, typically from an individual entering or exiting it. Thus, one possible way to cheat this method of motion detection is to move extremely slow so that eigenvalue spikes remain below the threshold. Despite this potential shortcoming, our work herein has shown that DMD can be applied to motion detection to produce a real-time, flexible, and efficient method that can also be used to isolate foreground motion once it has been detected.
Acknowledgements.
@@
Conflict of interest
The authors declare that they have no conflict of interest.

### v2

@@ 3 Motion Detection with DMD
Having reviewed the necessary DMD-related methods, we now present our method of motion detection. This application of DMD is the main contribution of this paper and the validation of the method is left to the next section. Recall that a visual overview of the method was presented in Figure 2.
@@
Input: $W\in\mathbb{R}^{M\times(N+1)}$, $r$, $p$, $T$
@@
The method starts by applying cDMD to windowed subsets of the streaming video data, which has been transformed into greyscale and vectorized into a matrix $W\in\mathbb{R}^{M\times(N+1)}$. The compression matrix $C$ and target rank $r$ are held constant over all windows, and each window is comprised of $T>0$ consecutive frames of video. Algorithm 1 provides our windowed cDMD method where we see that the output is the modulus of the continuous-time eigenvalues (5) within each window. The scope of this work is primarily limited to video data where the camera position remains fixed, meaning the background is assumed to be relatively unchanged over the frames. This means that if there is little or no motion in the video, the eigenvalues within the window will all correspond approximately to the background, i.e., $\omega_{i}\approx 0$. However, if motion is present in a window one expects to see this represented as at least one of the eigenvalues deviating from the background cluster. As one sweeps across windows, motion entering the frames is expected to be represented by a large spike in the DMD matrix spectrum, as shown in Figure 6 below.
Motion detection from the DMD spectrum is automated by considering the sets of eigenvalues $\{\omega_{i}^{(k)}\}_{i=1}^{r}$ and $\{\omega_{i}^{(k+1)}\}_{i=1}^{r}$ from the consecutive $k$th and $(k+1)$th windows. We compute the averages
@@
Then, defining a threshold $\Delta^{*}>0$, motion is flagged in the $(k+1)$th window if
@@
By tracking the relative change in the mean of the eigenvalues over windows we allow for the method to exclude background motion that persists over frames, such as leaves being rustled by the wind. Note that we need only save information from the $k$th and $(k+1)$th consecutive windows at a time, while if no motion is detected we discard information from the $k$th window and compare the spectrum of the $(k+1)$th and $(k+2)$th windows. Combined with the data reductions from cDMD, this means that only two $r\times r$ matrices are stored in memory to perform the method.
In practice, it was observed that foreground motion in video was sometimes accompanied by multiple consecutive activations; that is, the moduli of the eigenvalues would continue to increase in a few of the windows immediately following that of the initial detection. We believe that this is related to slow motion in the video, but we do not have a clear explanation for it. One way this can be resolved is by storing the (binary) results of the motion detection test for the previous few windows and additionally checking that no motion was already flagged in these previous windows.
Finally, if motion is detected in a window, we apply the method of background separation from Section 2.4. This allows for the isolation of the detected motion, again using only information from the stored $r\times r$ DMD matrix corresponding to the window where motion is first detected. This is summarized in the flowchart in Figure 4.
@@
This method involves initializing a number of parameters that may affect its performance. Specifically, the target rank $r$ dictates the number of eigenvalues in each window, while the target dimension $p$ from cDMD, the window length $T$, and most importantly, the detection threshold $\Delta^{*}$ have an impact on performance. The choice of parameters that leads to optimal performance of the method is problem-dependent and should be optimized with simple tests. This might include designing motion to be captured by the method, such as having someone walk into the frame, while also decreasing the sensitivity to exclude frivolous motion that might be unique to the application. In the following section we demonstrate the method’s performance on real video data, while also evaluating performance with receiver operating characteristic curves and tuning the threshold detection parameter with $k$-fold cross-validation.
@@ 4 Experimental Validation
In this section we demonstrate the performance of our method on real video data. A full code repository is available to reproduce the results of this section. This includes all files used for our tests, as well as full MATLAB scripts to perform the method on video data and all the following validations: https://github.com/marco-mig/dmd-motion-detection.
In what follows, we fix the parameters $T$, $p$, and $r$, while focusing on the importance of the threshold parameter $\Delta^{*}$. Specifically, we take $p=20$ and $r=5$, which means that each DMD matrix is only $5\times 5$. Thus, eigenvalues within each window can be computed quickly without presenting a bottleneck to the real-time nature of the algorithm. The window sizes are fixed to be $T=80$, corresponding to 8/3 seconds. The random matrix $C$ is constructed only once and applied to all videos for consistency.
@@ 4.1 Dataset
Initial tests of the method featured free videos from the Background Models Challenge database [50], including the one featured in Figure 1. However, to ensure a diversity of situations across the test set, we have opted to film our own videos using the built-in camera on the iPhone 13. Our videos were filmed to simulate security footage in that the camera is held still and various individuals enter and exit the frame to be detected. These videos include numerous variations such as the speed of entry and/or exit of the individuals, the distance of the motion from the camera, the time the individuals spend in the frame, and the actions performed by the individuals in the frame. Our test set includes a total of 20 videos, of which 15 were filmed in daylight and 5 with indoor lighting. Videos are of size 426 by 240 pixels and range in length from 27 seconds (810 frames) to 109 seconds (3270 frames). Each video begins and ends with a few seconds of a still background shot to provide a baseline for the method to gauge performance. Frames from videos in two lighting conditions are provided in Figure 5.
We use this video database to gauge the performance of our method and discuss optimal parameter tuning. In Section 4.5 we will then turn to the Microsoft Wallflower dataset [47] to provide further performance tests on a benchmark data set. We emphasize that for each new setting or camera placement one will be required to re-tune the parameters since the behavior of our method of motion detection depends on the specifics of the surrounding area. This means that one should not apply this method on streaming video footage without first optimally tuning the parameters on test footage from the same camera. This will be more apparent in Section 4.5 and, in particular, Table 2, where one sees that different threshold parameters are required to detect motion in the Microsoft Wallflower dataset due to the specific characteristics of each video, such as latent motion in the background.
@@ 4.2 Performance Demonstration
Our experiments demonstrate that sudden motion from an individual entering or exiting the video frame is accompanied by a spike in the average eigenvalues, as measured by (16). A characteristic demonstration is provided in Figure 6 where one can clearly see motion detection through large deviations in the DMD spectrum. When an individual enters the frame, it is reflected in an eigenvalue spike corresponding to the frame number minus the window length $T$. This is when the frame at which the individual enters the video first appears in the window as the rightmost column of the data matrix $W$, using the notation of Algorithm 1. Alternatively, when an individual exits the video, the window number where the eigenvalue spike takes place corresponds to the frame number itself. This is because it is the last window in which information about motion is present in the data matrix $W$, and so the change occurs after this information (in the left column of the data matrix) vanishes in the subsequent window. In summary, there is a temporal delay of length $T$, the window size, in detecting when an object has left the frame. Thus, a distinction is made between objects entering and exiting the frame, and the frame numbers supplied to the algorithm that calculates the error score in the following subsection have been adjusted accordingly.
@@ 4.3 Performance Evaluation
To display the relationship between the true positive rate (TPR) and false positive rate (FPR) we will use a receiver operating characteristic (ROC) curve (see, e.g., [21]). The TPR can be thought of the rate at which the algorithm successfully classifies positive events as positive, while the FPR is the rate at which the algorithm classifies negative events as positive. A perfect classifier necessarily has a TPR of 1 and an FPR of 0, and the ROC curve for such a classifier is the graph of the indicator function of the interval $(0,1]$. Our ROC curve is traced out by recording the TPR and FPR of the algorithm over threshold values $\Delta^{*}$ ranging from $10^{-10}$ to $10^{10}$. In this section, we do not implement the fix for consecutive activations as described in Section 3 as this interferes with the expected behaviour of the algorithm at threshold values close to zero. That is, for these values, activations at nearly every (if not every) window containing motion are expected. Finally, a half-second tolerance (forward and backward) in detecting each event is imposed.
@@ 4.4 Optimizing the Threshold
While our method contains a number of parameters to be set by the user, none is as important as the threshold $\Delta^{*}$. In this subsection we present a method for determining an optimal threshold to get the best performance from the method using a database of test videos that should be curated by the user before applying the method for video surveillance. We begin by defining an appropriate error measure which can then be minimized by the optimal threshold.
@@ 4.4.1 Measuring Error
Recall from Section 4.3 that there are a total of four possible outcomes of the algorithm: either it classifies correctly (encompassing TPs and TNs) or incorrectly (FPs and FNs). In many applications, these two latter outcomes are not equally costly. This is because in the context of video surveillance, a FN, such as not detecting an intruder, can be much more costly than a FP, such as falsely indicating the presence of an intruder. Thus, our goal here is to present a possible error metric that takes this into consideration.
@@ 4.4.2 A Variant of $k$-Fold Cross-Validation
We will use a variant of $k$-fold cross-validation (see, e.g., [21]) to determine an optimal threshold value using our error metric (17) and database of videos. In what follows we fix $d^{*}=30$, which corresponds to a tolerance of detection within 1 second of an event taking place. The penalty ratio in the error metric is taken to be $c=100$.
@@
Error $E_{i}$ 67.8 72.2 49.0 170.8
We implement the above with $k=4$ since our dataset is relatively small, and we initialize $P$ with 1000 evenly-spaced values between 0 and 1. Table 1 presents our results for one application of the modified $k$-fold cross-validation procedure. It can be seen that the optimal threshold corresponds to an error score of 49, where we recall that FP and FN have a score of 1 and 100, respectively. However, due to the small size of the video databases, there are large discrepancies between the error scores corresponding to the optimal thresholds. Moreover, each possible arrangement of the database into the training and validation sets results in different values. We expect that this cross-validation method will produce a more representative set of optimal thresholds and corresponding errors when applied to a larger database.
@@
From these results, two observations can be made. First, while the eigenvalue detection system can be effective when optimized for a single video, as shown in Figure 6, when optimizing over a set of videos, the performance tends to decrease. In particular, more FPs are made in order to minimize the occurrence of FNs over all the videos in the training set. This is evidently a consequence of the weight given to FN errors, and is thus dependent on the application. The second observation is that the optimal thresholds and their corresponding errors are always a trade-off between FPs and FNs, since they vary heavily depending on how each type of error is weighed according to the value of $c$. As demonstrated in Figure 8, plotting the average error over each of the training sets to which the cross-validation procedure is applied results in a curve with a distinct global minimum. Large error is observed for very small and very large threshold values due to the abundance of FP and FN errors, respectively.
4.5 Application to Benchmark Video Database
The previous demonstrations were carried out on a carefully curated database of videos that sought to probe performance based on variations in lighting, movement speed, and the number of objects moving. In this subsection we further demonstrate the performance of the method by implementing it on a benchmark set of publicly available videos. We use the Microsoft Wallflower dataset [47], which was recently featured in [37] to test object tracking with a Gaussian mixture model and in [23] to test a method of background modeling under variations in illumination across the frames. We further comment that tests were also performed on the Background Models Challenge Database [50], but results were felt to be unsatisfying for presentation herein since there is little movement into and out of the videos to detect with our method. Nonetheless, results of background separation with DMD were presented previously in Figure 1.
The Microsoft Wallflower dataset contains seven videos that are briefly described as follows.
  • Camouflage (CAM): A scene of a computer monitor with interference bars; a man enters and exits the scene.
  • ForegroundAperture (FA): A person is sleeping at a desk. They arise and exit the frame, then return and remain motionless at the desk before leaving again.
  • LightSwitch (LS): A computer lab scene; a person enters the room, turns the light off and leaves; another person enters, turns on the light, and remains for a while before leaving.
  • MovedObject (MO): A conference room scene; a person enters and pulls up to a chair to use the telephone, leaves the scene, and returns to use the chair and phone again, leaving them in a different position each time.
  • TimeOfDay (TOD): A room with couches gradually illuminates from darkness, a person enters and exits twice to sit on the couch; the scene then gradually returns to darkness.
  • WavingTree (WT): An outdoor scene in daylight with a tree swaying in the background; a person briefly enters and exits the scene.
  • Bootstrap (BS): An overhead view of a busy cafeteria scene with many people entering and exiting the camera’s view. Due to the presence of many subjects moving into and out of the frame constantly, our algorithm is unable to discern any foreground motion.
We reiterate that, due to the nature of sliding window DMD, events in the first $T$ and last $T$ frames of the video are outside the scope of detection since there is nothing to compare the DMD spectrum with yet. Thus, any event(s) in the database videos that fall into this regime will not be incorporated into the error score calculation. Table 2 provides both the optimal threshold value for each of the above videos and the corresponding error score. All other parameter values are fixed across all videos in this database and are the same as in the previous tests from our own curated database.
Video name: CAM FA LS MO TOD WT BS
Error Score 1 16 24 1 15 2 N/A
Optimal Threshold 0.1530 0.5020 0.3850 0.5400 0.5010 0.1850 N/A
Table 2: Result of applying the motion detection algorithm to the Microsoft Wallflower database.
We can see from Table 2 that our method performs well on three of the videos (CAM, MO, and WT), with only one or two false positives detected. The strong performance on WT and CAM in particular is an improvement over the results in [37], where their method displayed some difficulties with these videos. Alternatively, we produce more false positives on the remaining three videos (FA, LS, and TOD). Two out of the three latter videos (LS and TOD) feature more pronounced lighting changes than the other videos, which could explain the algorithm’s difficulty in detecting foreground motion since this is a well-known issue for many motion detection algorithms (see [23] and the references therein). Overall, we can see that the optimal threshold varies widely for each video. This emphasizes the importance of recalibrating the algorithm for every new setting, as the same threshold value may not work well across varied settings.
@@ 5 Conclusion
The method relies on spikes in the DMD spectrum coming from sudden motion in the frame, typically from an individual entering or exiting it. The height of the spikes is likely to be related to the speed of movement entering the frame, since larger DMD eigenvalues are required to produce faster movement in the DMD decomposition using the temporal components $\mathrm{e}^{\omega_{m}t}$. Thus, one possible way to fool this method of motion detection is to move extremely slowly so that eigenvalue spikes remain below the threshold. Furthermore, the spikes in the DMD spectrum related to movement are often short-lived (see Figure 6) and so it is unclear if their duration indicates anything important about other movement factors such as volume and number of objects. Despite these potential shortcomings, our work herein has shown that DMD can be applied to motion detection to produce a real-time, flexible, and efficient method that can also be used to isolate foreground motion once it has been detected.
Availability of Data and Material
All code to reproduce the demonstrations in this manuscript is available in the repository https://github.com/marco-mig/dmd-motion-detection. The demonstrations in Section 4.5 use videos from the Microsoft Wallflower database, available at https://www.microsoft.com/en-us/download/details.aspx?id=54651. Our own video database will be made available to interested users upon reasonable request.
Competing Interests
The authors declare that they have no competing interests.
Funding
This research was supported by an NSERC USRA (MM) and NSERC Discovery Grants (SB and JJB).
Authors’ Contributions
All authors contributed to the study conception and design. Numerical experiments were performed by MM, under the supervision of SB and JJB. All authors contributed equally to drafting the previous versions of the manuscript, as well as reading and approving the final manuscript. MM created all figures.
Acknowledgements